# A new general quantum state verification protocol by the classical shadow method

### v3

@@ Abstract
Verifying whether a quantum device produces a specific state is a fundamental task in many applications of modern quantum technologies. In the conventional framework of quantum state verification, designing an optimal or efficient protocol for each type of state often requires intricate customization. Recently, Hsin-Yuan Huang et al. introduced a novel approach called the shadow overlap protocol [1], which leverages classical shadows to efficiently verify multiple types of states simultaneously. In this paper, we propose a new verification protocol that integrates key ideas from both the conventional framework and the shadow overlap protocol. To achieve this, we first reformulate the shadow overlap protocol using the language of hypothesis testing, which also underpins the conventional framework, and then analyze the similarities and differences between them. Our approach enhances the capabilities of the shadow overlap protocol while addressing some of its limitations, leading to improved sample complexity and a more natural treatment of states with specific structures. Finally, we demonstrate the effectiveness of our protocol by applying it to GHZ states and stabilizer states.
@@ I Introduction
The traditional quantum state tomography methods can’t be used to solve the QSV problem, as they need exponential copies of the prepared states with respect to the number of qubits to obtain a classical description of the state. To circumvent such a predicament, Sam Pallister et al. proposed a general framework, called the PLM framework, for verifying the prepared quantum states in [2]. An explicit verification strategy in this framework consists of a sequence of tests constructed by local projective measurements with binary outcomes, if the prepared states pass all such tests, we will accept the statement that the device produces the target state correctly. The sample complexity, i.e., the number of prepared states, of such a general framework is approximately proportional to the inverse of the spectral gap of an operator corresponding to the strategy, called the strategy operator. Hence, applying this framework to verify a specific type of state reduces to constructing an explicit protocol or a strategy operator, which has the maximal spectral gap.
In fact, before the proposition of the PLM framework, there have been some related works, like [3, 4] trying to verify the maximally entangled states. After the PLM framework, there are also many developments in the direction of QSV. The optimal protocol (strategy operator) with the maximal spectral gap has been constructed for states, like the Bell states, the maximally entangled states, the bipartite pure state, the GHZ states and the stabilizer states [2, 5, 6, 7, 8], and some near-optimal or efficient protocols are also proposed for some more complicated states like the W states, the Dick states, the phased Dick states, graph states and hypergraph states [9, 10, 11, 12, 13]. In addition to the mentioned protocols, there are also some works trying to extend the PLM framework to more complicated or more practical circumstances, like the LOCC measurement [6, 14, 9, 15, 10], the adversarial scenario [16, 17], the blind measurement-based quantum computation [18].
@@
Fortunately, Hsin-Yuan Huang et al. proposed a new, different protocol in [1], which can verify almost all Haar random pure quantum states and some special states efficiently. This protocol, called the shadow overlap protocol (SOP), unifies the ideas of randomized Pauli measurement and the classical shadow tomography [19, 20], then possesses the advantages of application in practical experiments. However, the SOP also has certain limitations. First, for some structured states such as GHZ states, applying the SOP requires additional unitary transformations, which are not inherently integrated into the protocol and must be handled separately. Second, the protocol relies on a complicated operator constructed by querying the amplitudes of the target state in the computational basis, which lacks a clear physical interpretation.
In this paper, we explore the similarities and differences between the SOP and the PLM framework, and try to remedy the shortcomings of the SOP by combining the ideas of both frameworks into a new protocol. More explicitly, we interpret the SOP in the view of hypothesis test and compare it with the PLM framework. Then we propose a new protocol, so-called the directly partial shadow overlap (DPSO) protocol, in the same framework of the SOP, which doesn’t need the queries of the target state. Compared with the SOP, our protocol has lower sample complexity and is more powerful, which can deal with the GHZ state or other special states more naturally. When the level of our protocol is equal to $1$, the level-$1$ DPSO protocol includes the SOP, so our protocol also retains the capability to verify almost all Haar random pure states.
@@ II Preliminaries
In this section, we will very briefly introduce the PLM framework of QSV proposed in [2] and its subsequent developments. The task of quantum state verification is to determine if a quantum device produces a particular target state accurately, and a verification protocol provides us an explicit method to assess whether the output of this device is the target state rather than other states. More explicitly, given a target state $|\psi\rangle$, a quantum device produces a sequence of states $\rho_{1},\cdots,\rho_{N}$, then we need to accept one of the following statements: (1) the device produces the target state exactly, i.e., $\rho_{i}=|\psi\rangle\langle\psi|$ for all $1\leq i\leq N$, or (2) the device can’t produce the target state exactly, meaning $\langle\psi|\rho_{i}|\psi\rangle\leq 1-\epsilon$ with $10<\epsilon<$ 1 for all $i$.
Figure 1: The illustration of the procedure of the PLM framework. A quantum device is accepted until it passes $N$ tests successively.
@@ II.1 The PLM framework of QSV
[…]
@@
[…]
@@ II.2 QSV with LOCC measurements
There have also been many improvements on the PLM framework of QSV, amongst which is the substitution of the local measurements by the measurements composed of local operations and classical communications (LOCC) [6, 14, 9, 15, 10]. In [2], the test operator $E_{i}$’s are constructed by local projective measurements of each party of the system and there are no classical communications among different parties. If the classical communications among the parties are allowed, there are more types of test operators can be constructed. Taking the one-way LOCC strategy operator $\Omega^{\rightarrow}$ of a bipartite system as an example, Alice performs a local measurement $\{M_{a|j}\}$ with probability $p_{j}$ on her subsystem, then Alice sends $j$ and the measurement outcome $a$ to Bob, and then Bob makes the measurement $\{K_{a|j},\mathds{1}-K_{a|j}\}$ on his subsystem based on $j,a$. So $\Omega^{\rightarrow}$ has the following form,
@@
The above procedure can also be generalized to two-way LOCC and many-round two-way LOCC. One special example of one-way LOCC strategy operators is proposed in [23]. In [23], the verifier is given a $n$-qubit state produced by a quantum device and apply a randomized local Pauli measurements to $n-1$ qubits randomly. After obtaining the measurement outcomes $\boldsymbol{s}=(s_{1},\cdots,s_{n-1})$, the verifier applies a projective measurement to the left qubit, where the projective measurement is just the reduced state of $|\psi\rangle$ with respect to the outcome $\boldsymbol{s}$. Astonishingly, such simple protocol is proved to be valid for many types of states numerically or analytically in [23].
@@ III.1 The shadow overlap protocol
Parallel to advancements in the PLM framework for QSV, Hsin-Yuan Huang et al. recently proposed a new general protocol termed shadow overlap protocol (SOP) in [1], which certifies nearly all Haar random pure states. This protocol utilizes the classical shadow method [20] and employs a classical query model $\mathcal{Q}(\psi)$, which is a classical description of the target state $|\psi\rangle$ and stored in a classical computer. At first glance, the SOP appears very different from the PLM framework, but it also shares some similarities with the PLM framework, particularly with the previous mentioned protocol in [23]. Here, we briefly review the SOP and reformulate it by using the terminologies of hypothesis testing to unify its description and the PLM framework and facilitate comparison.
Figure 2: Illustration of the SOP procedure for $n=4,\ell=1$. The left blue rectangle represents the $4$-qubit state $\rho$ produced by a quantum device, while the right dashed rectangle represents the query model of the target state $|\psi\rangle$ stored in a classical computer. On the left side, the $2$nd qubit is sampled in the first step and Pauli-$Z$ measurements are applied to the remaining three qubits, yielding measurement outcomes $(z_{1},z_{3},z_{4})$. The classical shadow of the $2$nd qubit’s state, $\hat{\sigma}_{(z_{1},z_{3},z_{4})}$, is then constructed from the outcome of the randomized Pauli measurement. On the right side, given the measurement outcomes $(z_{1},z_{3},z_{4})$, the classical model is queried to obtain the amplitudes corresponding to $\tilde{\boldsymbol{z}}_{1}=(z_{1},0,z_{3},z_{4})$ and $\tilde{\boldsymbol{z}}_{2}=(z_{1},1,z_{3},z_{4})$. After constructing the projector $L_{(z_{1},z_{3},z_{4})}$, the shadow overlap is computed classically.
Suppose the $n$-qubit target state is $|\psi\rangle$ with its classical query model $\mathcal{Q}(\psi)$ stored in a classical computer and a quantum device produces a state $\rho$, the procedure of each run of the $-\ell$ SOP comprises of two stages.
  • The first stage mainly consists of the measurements on quantum state $\rho$ as following:
@@
      Perform Pauli-$Z$ measurements to the left $n-r$ qubits, obtain the outcomes $\boldsymbol{z}_{\bar{K}}=(z_{1},\cdots,z_{n-r})\in\{0,1\}^{n-r}$ with $\bar{K}$ being the complement of $K$.
@@ III.2 The hypothesis testing
First, we choose the same null and alternative hypotheses as those in the PLM framework. Before considering two type of errors, we need to select a decision rule, which determines which hypothesis is accepted according to the value of estimator $\hat{\omega}_{\text{a}}$ [21, 22]. If $H_{1}$ is true, $\mathop{\rm Tr}(L\rho)=1$, then as an estimator of $\mathop{\rm Tr}(L\rho)$, $\hat{\omega}_{a}$ fluctuates around $1$. While if $H_{0}$ is true, $\mathop{\rm Tr}(L\rho)\leq 1-\nu(L)\epsilon$ with $\nu(L)$ being the spectral gap of $L$, then $\hat{\omega}_{a}$ fluctuates around $\mathop{\rm Tr}(L\rho)$ but is less than $1-\nu(L)\epsilon$ probably. Hence, we tend to reject $H_{0}$ and accept $H_{1}$ if $\hat{\omega}_{a}>1-\nu(L)\epsilon$. So we set the decision rule as:
@@
with $11-\nu(L)\epsilon<t_{0}<$ 1. In fact, different values of $t_{0}$ generate different tests, but it will be much simpler when setting $t_{0}=1-\nu(L)\epsilon/2$ as will be seen.
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Returning to the discussion of the SOP, the decision rule is: if the estimated shadow overlap $\hat{\omega}_{\text{e}}\geq 1-\epsilon\nu(L)/2$, then we conclude that the quantum device produces the target state. According to Eq.(10), the value of $\hat{\omega}$ is bounded as $0<\hat{\omega}<2^{2\ell-1}$ as proved in [1], then we get the explicit sample complexity
@@ III.3 Comparison with the PLM framework
There are many striking similarities between the SOP and the PLM framework. First, the operator $L$ defined in Eq.(10) has the same properties as the strategy operator $\Omega$ in the PLM framework: $0\leq L\leq\mathds{1}$ and $L|\psi\rangle=|\psi\rangle$. Second, the procedure of randomized measurement in the SOP is very similar with that in [23], especially the case of $\ell=1$.
However, there are also some fundamental differences between the SOP and the PLM framework. First, the measurement applied in each run of the SOP is not a LOCC measurement but rather a local Pauli measurement, as there is no classical communication among $n$ parties before performing the Pauli measurements. This fundamental distinction sets it apart from the LOCC measurements described in Eq.(5). Although the Pauli measurements are performed on $n$ qubits in two steps during the first stage of the SOP, in a real experiment, they can instead be executed simultaneously, with the classical shadow computed in the second stage. Second, although we have reformulated the SOP using hypothesis testing terminology, the resulting form differs from standard hypothesis testing as described in [21, 22]. The SOP more closely resembles parameter estimation, as its objective is to estimate $\mathop{\rm Tr}(L\rho)$. In the PLM framework, $\mathop{\rm Tr}(\rho\Omega)$ is the probability for $\rho$ passing a test and the decision rule is determined by the number of ”pass”. While in the SOP, we need to estimate $\mathop{\rm Tr}(\rho L)$ and make a decision based on the estimated value. Thirdly, as shown in Eq.(24), the sample complexity of the SOP appears significantly larger than that of the PLM framework due to the exponential term $2^{4\ell-1}$. However, when $\ell$ is set to a small value, this term becomes negligible compared to other factors. Regarding other factors, the sample complexity $N$ of the SOP scales as $(\epsilon^{2}\nu^{2}(L))$, which is the consequence of the Hoeffding’s inequality. While this scaling is less favorable, it enables the SOP to certify a broader class of states. In other words, there is a trade-off between the capability of uniformly verifying different types of state and the sample complexity of the protocol.

### v4

@@ Abstract
Verifying whether a quantum device produces a specific quantum state is a fundamental task in many applications of modern quantum technologies. In the conventional framework of quantum state verification, designing an optimal or efficient protocol for each type of state typically requires intricate, state-specific customization. Recently, Hsin-Yuan Huang et al. introduced a novel approach known as the shadow overlap protocol Huang et al. (2024a), which leverages classical shadows to efficiently verify multiple classes of quantum states simultaneously. In this work, we propose a new verification protocol that integrates key ideas from both the conventional framework and the shadow overlap protocol. To this end, we first reformulate the shadow overlap protocol using the formalism of hypothesis testing, which also underpins the conventional approach, and analyze the similarities and differences between the two. Our framework extends the capabilities of the shadow overlap protocol while addressing some of its limitations, yielding improved sample complexity and a more natural treatment of structured quantum states. We demonstrate the effectiveness of our protocol through applications to GHZ states and stabilizer states.
Quantum state verification, classical shadow, Pauli measurements, Stabilizer states, GHZ states
@@ I Introduction
The traditional quantum state tomography methods can’t be used to solve the QSV problem, as they need exponential copies of the prepared states with respect to the number of qubits to obtain a classical description of the state. To circumvent such a predicament, Sam Pallister et al. proposed a general framework, called the PLM framework, for verifying the prepared quantum states in Pallister et al. (2018). An explicit verification strategy in this framework consists of a sequence of tests constructed by local projective measurements with binary outcomes, if the prepared states pass all such tests, we will accept the statement that the device produces the target state correctly. The sample complexity, i.e., the number of prepared states, of such a general framework is approximately proportional to the inverse of the spectral gap of an operator corresponding to the strategy, called the strategy operator. Hence, applying this framework to verify a specific type of state reduces to constructing an explicit protocol or a strategy operator, which has the maximal spectral gap.
Prior to the introduction of the PLM framework, several related works Hayashi et al. (2006); Hayashi (2009) had already explored the verification of maximally entangled states. Notably, the authors of Hayashi (2009) first addressed the quantum state verification under the assumption that the states produced by the quantum device are independently and identically distributed (i.i.d.). In contrast, the PLM framework considers a more general setting in which the states may be adversarial and are therefore not necessarily i.i.d. After the PLM framework, there are also many developments in the direction of QSV. The optimal protocol (strategy operator) with the maximal spectral gap has been constructed for states, like the Bell states, the maximally entangled states, the bipartite pure state, the GHZ states and the stabilizer states Pallister et al. (2018); Zhu and Hayashi (2019a); Yu et al. (2019); Li et al. (2020); Dangniam et al. (2020), and some near-optimal or efficient protocols are also proposed for some more complicated states like the W states, the Dick states, the phased Dick states, graph states and hypergraph states Liu et al. (2019); Li et al. (2021); Zhu and Hayashi (2019b); Chen et al. (2023); Zhu et al. (2024). In addition to the mentioned protocols, there are also some works trying to extend the PLM framework to more complicated or more practical circumstances, like the LOCC measurement Yu et al. (2019); Wang and Hayashi (2019); Liu et al. (2019); Yu et al. (2022); Li et al. (2021), the adversarial scenario Zhu and Hayashi (2019c, d), the blind measurement-based quantum computation Li et al. (2023).
@@
Fortunately, Hsin-Yuan Huang et al. proposed a new, different protocol in Huang et al. (2024a), which can verify almost all Haar random pure quantum states and some special states efficiently. This protocol, called the shadow overlap protocol (SOP), unifies the ideas of randomized Pauli measurement and the classical shadow tomography Elben et al. (2022); Huang et al. (2020), then possesses the advantages of application in practical experiments. However, the SOP also has certain limitations. First, for some structured states such as GHZ states, applying the SOP requires additional unitary transformations, which are not inherently integrated into the protocol and must be handled separately. Second, the protocol relies on a complicated operator constructed by querying the amplitudes of the target state in the computational basis, which lacks a clear physical interpretation.
In this paper, we explore the similarities and differences between the SOP and the PLM framework, and propose a new protocol, so-called the directly partial shadow overlap (DPSO) protocol, which remedies some shortcomings of the SOP by combining the ideas of both frameworks. More explicitly, we reinterpret the SOP in the view of hypothesis test and compare it with the PLM framework. Then we construct the DPSO protocol within the same framework as SOP, under the assumption that the states generated by the quantum device are i.i.d. Compared with the SOP, our protocol has lower sample complexity and broader applicability, being able to handle GHZ states or other special states more naturally. When the level of our protocol is equal to $1$, the level-$1$ DPSO protocol includes the SOP, so our protocol also retains the capability to verify almost all Haar random pure states.
@@ II Preliminaries
In this section, we will very briefly introduce the PLM framework of QSV proposed in Pallister et al. (2018) and its subsequent developments. The task of quantum state verification is to determine if a quantum device produces a particular target state accurately, and a verification protocol provides us an explicit method to assess whether the output of this device is the target state rather than other states. More explicitly, given a target state $|\psi\rangle$, a quantum device produces a sequence of states $\rho_{1},\cdots,\rho_{N}$, which may be i.i.d. or adversarial, then we need to accept one of the following statements: (1) the device produces the target state exactly, i.e., $\rho_{i}=|\psi\rangle\langle\psi|$ for all $1\leq i\leq N$, or (2) the device can’t produce the target state exactly, meaning $\langle\psi|\rho_{i}|\psi\rangle\leq 1-\epsilon$ with $0<\epsilon<1$ for all $i$.
Figure 1: The illustration of the procedure of the PLM framework. A quantum device is accepted until it passes NN tests successively.
@@ II.1 The PLM framework of QSV
$\text{Pr}[\text{"Pass"}]=\sum_{j}p_{j}\mathop{\rm Tr}(E_{j}\rho_{i})=\mathop{\rm Tr}(\Omega\rho_{i}),$ (1)
@@
On the assumption of $E_{j}|\psi\rangle=|\psi\rangle$, the Type II error, which is the probability of accepting $H_{0}$ given $H_{1}$ is true, is always 0. Therefore, there is a unique optimal acceptance set or decision rule in the view of the classical hypothesis testing Cover and Thomas (2006); Hogg et al. (2013), which is just the rule shown in Protocol 1. The Type I error, the probability of accepting $H_{1}$ given $H_{0}$ is true, is just the probability of all $N$ states $\rho_{1},\cdots,\rho_{N}$ passing the test,
$\text{Pr}[\text{"Accept"}]=\prod_{i=1}^{N}\mathop{\rm Tr}(\Omega\rho_{i})\leq[1-\nu(\Omega)\epsilon]^{N}.$ (3)
Then by requiring the probability $\text{Pr}[\text{"Accept"}]$ being smaller than a positive constant $\delta$, we can get the lower bound of the sample complexity,
@@ II.2 QSV with LOCC measurements
There have also been many improvements on the PLM framework of QSV, amongst which is the substitution of the local measurements by the measurements composed of local operations and classical communications (LOCC) Yu et al. (2019); Wang and Hayashi (2019); Liu et al. (2019); Yu et al. (2022); Li et al. (2021). In Pallister et al. (2018), the test operator $E_{i}$’s are constructed by local projective measurements of each party of the system and there are no classical communications among different parties. If the classical communications among the parties are allowed, there are more types of test operators can be constructed. Taking the one-way LOCC strategy operator $\Omega^{\rightarrow}$ of a bipartite system as an example, Alice performs a local measurement $\{M_{a|j}\}$ with probability $p_{j}$ on her subsystem, then Alice sends $j$ and the measurement outcome $a$ to Bob, and then Bob makes the measurement $\{K_{a|j},\mathds{1}-K_{a|j}\}$ on his subsystem based on $j,a$. So $\Omega^{\rightarrow}$ has the following form,
@@
The above procedure can also be generalized to two-way LOCC and many-round two-way LOCC. One special example of one-way LOCC strategy operators is proposed in Li and Zhu (2025). In Li and Zhu (2025), the verifier is given a $n$-qubit state produced by a quantum device and apply a randomized local Pauli measurements to n1n-1 qubits randomly. After obtaining the measurement outcomes $\boldsymbol{s}=(s_{1},\cdots,s_{n-1})$, the verifier applies a projective measurement to the remaining qubit, where the projective measurement is just the reduced state of $|\psi\rangle$ with respect to the outcome $\boldsymbol{s}$. Astonishingly, such simple protocol is proved to be valid for many types of states numerically or analytically in Li and Zhu (2025).
@@ III.1 The shadow overlap protocol
Parallel to advancements in the PLM framework for QSV, Hsin-Yuan Huang et al. recently proposed a new general protocol termed shadow overlap protocol (SOP) in Huang et al. (2024a), which certifies nearly all Haar random pure states. This protocol utilizes the classical shadow method Huang et al. (2020) and employs a classical query model $\mathcal{Q}(\psi)$, which is a classical description of the target state $|\psi\rangle$ and stored in a classical computer. At first glance, the SOP appears very different from the PLM framework, but it also shares many similarities with the PLM framework, particularly with the previous mentioned protocol in Li and Zhu (2025). Note that the SOP is designed to handle only i.i.d. states produced by the quantum device. Here, we briefly review the SOP and reformulate it by using the terminologies of hypothesis testing to unify its description and the PLM framework and facilitate comparison.
Figure 2: Illustration of the SOP procedure for $n=4,\ell=1$. The left blue rectangle represents the $4$-qubit state $\rho$ produced by a quantum device, while the right dashed rectangle represents the query model of the target state $|\psi\rangle$ stored in a classical computer. On the left side, the 22nd qubit is sampled in the first step and Pauli-$Z$ measurements are applied to the remaining three qubits, yielding measurement outcomes $(z_{1},z_{3},z_{4})$. The classical shadow of the 22nd qubit’s state, $\hat{\sigma}_{(z_{1},z_{3},z_{4})}$, is then constructed from the outcome of the randomized Pauli measurement. On the right side, given the measurement outcomes $(z_{1},z_{3},z_{4})$, the classical model is queried to obtain the amplitudes corresponding to $\tilde{\boldsymbol{z}}_{1}=(z_{1},0,z_{3},z_{4})$ and $\tilde{\boldsymbol{z}}_{2}=(z_{1},1,z_{3},z_{4})$. After constructing the projector $L_{(z_{1},z_{3},z_{4})}$, the shadow overlap is computed classically.
Suppose the $n$-qubit target state is $|\psi\rangle$ with its classical query model $\mathcal{Q}(\psi)$ stored in a classical computer and a quantum device produces states $\rho$’s which are i.i.d., the procedure of each run of the level-$\ell$ SOP comprises of two stages.
  • The first stage mainly consists of the measurements on a quantum state $\rho$ as following:
@@
      Perform Pauli-$Z$ measurements to the remaining nrn-r qubits, obtain the outcomes $\boldsymbol{z}_{\bar{K}}=(z_{1},\cdots,z_{n-r})\in\{0,1\}^{n-r}$ with $\bar{K}$ being the complement of $K$.
@@ III.2 The hypothesis testing
First, we choose the same null and alternative hypotheses as those in the PLM framework. Before considering two type of errors, we need to select a decision rule, which determines which hypothesis is accepted according to the value of estimator $\hat{\omega}_{\text{a}}$ Cover and Thomas (2006); Hogg et al. (2013). If $H_{1}$ is true, $\mathop{\rm Tr}(L\rho)=1$, then as an estimator of $\mathop{\rm Tr}(L\rho)$, $\hat{\omega}_{a}$ fluctuates around $1$. While if $H_{0}$ is true, $\mathop{\rm Tr}(L\rho)\leq 1-\nu(L)\epsilon$ with $\nu(L)$ being the spectral gap of $L$, then $\hat{\omega}_{a}$ fluctuates around $\mathop{\rm Tr}(L\rho)$ but is less than $1-\nu(L)\epsilon$ probably. Hence, we tend to reject $H_{0}$ and accept $H_{1}$ if $\hat{\omega}_{a}>1-\nu(L)\epsilon$. So we set the decision rule as:
@@
with $1-\nu(L)\epsilon<t_{0}<1$. In fact, different values of $t_{0}$ generate different tests, but it will be much simpler when setting $t_{0}=1-\nu(L)\epsilon/2$ as will be seen.
Next we try to compute the Type I and Type II errors. The Type I error is
@@
The upper bounds of Type I and Type II errors in Eq.(15) and Eq.(17) show that the speeds of both types of error approaching 0 are exponentially fast with respect to $N$. But if $N$ is fixed, when $t_{0}$ approaches $1$, the Type II error becomes larger and the Type I error becomes smaller. It’s just the reverse case when $t_{0}$ approaches $1-\nu(L)\epsilon$. There is a tradeoff between the two errors, so we have to let one error be smaller than a small positive constant $\chi$ and compute the minimum of another error, just as what we should do in the standard hypothesis testing Cover and Thomas (2006); Hogg et al. (2013). After requiring the second minimal error to be smaller than $\delta$, then we can compute the sample complexity as shown in the following theorem.
@@
Suppose that the Type II error $\beta:=\rm{Pr}[\rm{reject}~H_{1}|H_{1}]$ is smaller than $\chi$ with $0<\chi<1$, then the Type I error $\alpha:=\rm{Pr}[\rm{reject}~H_{0}|H_{0}]$ takes the optimal (minimal) value $\alpha^{*}$ when
@@
As $\alpha$ decreases when $t_{0}$ increases, so $\alpha$ is minimal when $t_{0}=1-(b-a)\sqrt{\ln\chi^{-1}/(2N)}$. Then we expect the Type I error being smaller than $\delta$, i.e.,
@@
In fact, we can also fix the Type I error $\alpha$ and compute the optimal Type II error $\beta$, but we get the same sample complexity as Eq.(20). If we require both types of error to be smaller than one small constant like $\delta$, the situation is much simplified, and $t_{0}$ takes the value $1-\nu(L)\epsilon/2$, then we can compute the sample complexity as
@@
Returning to the discussion of the SOP, the decision rule is: if the estimated shadow overlap $\hat{\omega}_{\text{e}}\geq 1-\epsilon\nu(L)/2$, then we conclude that the quantum device produces the target state. According to Eq.(10), the value of $\hat{\omega}$ is bounded as $0<\hat{\omega}<2^{2\ell-1}$ as proved in Huang et al. (2024a), then we get the explicit sample complexity
@@ III.3 Comparison with the PLM framework
There are many striking similarities between the SOP and the PLM framework. First, the operator $L$ defined in Eq.(10) has the same properties as the strategy operator $\Omega$ in the PLM framework: $0\leq L\leq\mathds{1}$ and $L|\psi\rangle=|\psi\rangle$. Second, the procedure of randomized measurement in the SOP is very similar with that in Li and Zhu (2025), especially the case of $\ell=1$.
However, there are also some fundamental differences between the SOP and the PLM framework. First, the SOP assumes that the quantum device produces the i.i.d. states, but the $\rho$’s can be adversarial in the PLM framework. Secondly, the measurement applied in each run of the SOP is not a LOCC measurement but rather a local Pauli measurement, as there is no classical communication among $n$ parties before performing the Pauli measurements. This fundamental distinction sets it apart from the LOCC measurements described in Eq.(5). Although the Pauli measurements are performed on $n$ qubits in two steps during the first stage of the SOP, in a real experiment, they can instead be executed simultaneously, with the classical shadow computed in the second stage. Thirdly, although we have reformulated the SOP using hypothesis testing terminology, the resulting form differs from standard hypothesis testing as described in Cover and Thomas (2006); Hogg et al. (2013). The SOP more closely resembles parameter estimation, as its objective is to estimate $\mathop{\rm Tr}(L\rho)$. In the PLM framework, $\mathop{\rm Tr}(\rho\Omega)$ is the probability for $\rho$ passing a test and the decision rule is determined by the number of ”pass”. While in the SOP, we need to estimate $\mathop{\rm Tr}(\rho L)$ and make a decision based on the estimated value. Fouthly, as shown in Eq.(24), the sample complexity of the SOP appears significantly larger than that of the PLM framework due to the exponential term $2^{4\ell-1}$. However, when $\ell$ is set to a small value, this term becomes negligible compared to other factors. Regarding other factors, the sample complexity $N$ of the SOP scales as $(\epsilon^{2}\nu^{2}(L))$, which is the consequence of the Hoeffding’s inequality. While this scaling is less favorable, it enables the SOP to certify a broader class of states. In other words, there is a trade-off between the capability of uniformly verifying different types of state and the sample complexity of the protocol.